OpenStudy (anonymous):

The manager of a furniture factory finds that it costs $2200 to manufacture 100 chairs in one day and $4000 to produce 300 chairs in one day. (a) Assuming that the relationship between cost C and the number of chairs produced x is linear, find an equation that expresses this relationship.

jimthompson5910 (jim_thompson5910):

You have two points (100,2200) and (300,4000) Find the equation of the line that goes through these points. Let me know if you need help doing that.

OpenStudy (anonymous):

i got y=9x-2600

jimthompson5910 (jim_thompson5910):

The slope is 9, but the y-intercept is NOT -2600

jimthompson5910 (jim_thompson5910):

y = mx+b y = 9x + b 2200 = 9(100) + b 2200 = 900 + b Keep going to solve for b

OpenStudy (anonymous):

b=1300 but i dont understand how you got that equation

jimthompson5910 (jim_thompson5910):

y = mx+b y = 9x + b ... Plug in m = 9 (the slope you found earlier) 2200 = 9(100) + b ... Plug in (100, 2200) or x = 100 and y = 2200 ....

OpenStudy (anonymous):

i see what i did wrong for some reason i used 100 and 300 instead of 2200 and 100

jimthompson5910 (jim_thompson5910):

Yes use the corresponding x (number) and y (cost) values

jimthompson5910 (jim_thompson5910):

x = 100 doesn't correspond to y = 300, but x = 100 corresponds to y = 2200

OpenStudy (anonymous):

i was still partially thinking slope instead of point-slope form

jimthompson5910 (jim_thompson5910):

i see

OpenStudy (anonymous):

the second and third parts of my question ask what my slope and y-int is and what they represent. I know what my slope and y-int is just not what they represent

jimthompson5910 (jim_thompson5910):

slope represents the change in cost over the change in number of chairs made

jimthompson5910 (jim_thompson5910):

so it's the cost per chair

jimthompson5910 (jim_thompson5910):

so it costs $9 a chair Notice how as x gets bigger by one, y will get bigger by 9 (because of the slope)

jimthompson5910 (jim_thompson5910):

The y-intercept is the starting cost to do business regardless of how many chairs are made (think of this as the cost of rent, electricity, etc etc)

jimthompson5910 (jim_thompson5910):

So it costs $1300 to keep the lights on and the machines running (no matter if 0 chairs or 100 chairs are made)

OpenStudy (anonymous):

Thank your so much for helping me!

jimthompson5910 (jim_thompson5910):

yw